1. Suppose that a seller decreases price for his/her good by 5 percent. Discuss how total revenue changes when : a) Ed=0.3  b) Ed=1 c) Ed=1.5


1
Expert's answer
2021-03-31T11:57:34-0400

When Ed=0.3, percentage change in quantity will slightly increase by "0.3\times 5\%=1.5\%" , so total revenue will increase by "0.3\times 5\%=1.5\%"

when Ed=1, total revenue will increase by "1\times5=5\%" when prices decrease by the said percentage.

when Ed=1.5, total revenue will increase by "1.5\times 5\%=7.5\%" since an elasticity of demand above 1 shows customers react more to changes in price.